Fifteen people took part in a free-throw competition. The first participant scored
x points, the second scored y points, the third scored the arithmetic mean of the
number of points scored by the first two, and each subsequent competitor scored
the arithmetic mean of the number of points scored by all previous ones. How
many points did the last competitor score?

(A) (x+y)/2 (B) 12(x+y)/14 (C) 13(x+y)/15
(D) 14(x+y)/16 (E) none of these

Answer by user_dude2008(1862)

(x+y+(x+y)/2)/3 = ((3x+3y)/2)/3 = (3x+3y)/6 = (x+y)/2

Answer: A)